# Gaussian Splashing: Dynamic Fluid Synthesis with Gaussian Splatting

### v1

@@ Abstract.
We demonstrate the feasibility of integrating physics-based animations of solids and fluids with 3D Gaussian Splatting (3DGS) to create novel effects in virtual scenes reconstructed using 3DGS. Leveraging the coherence of the Gaussian splatting and position-based dynamics (PBD) in the underlying representation, we manage rendering, view synthesis, and the dynamics of solids and fluids in a cohesive manner. Similar to Gaussian shader, we enhance each Gaussian kernel with an added normal, aligning the kernel’s orientation with the surface normal to refine the PBD simulation. This approach effectively eliminates spiky noises that arise from rotational deformation in solids. It also allows us to integrate physically based rendering to augment the dynamic surface reflections on fluids. Consequently, our framework is capable of realistically reproducing surface highlights on dynamic fluids and facilitating interactions between scene objects and fluids from new views. For more information, please visit our project page at https://amysteriouscat.github.io/GaussianSplashing/.
copyright: nonesubmissionid: 303
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 1. Splashing LEGO.  Gaussian Splashing (GSP) is a unified framework combining position-based dynamics and 3D Gaussian Splatting (3DGS). By leveraging their coherent point-based representations, GSP delivers high-quality rendering for novel dynamic views involving interacting solids and fluids. GSP enables a variety of interesting effects and new human-computer interaction modalities that are not available with existing NeRF/3DGS based systems. The teaser figure showcases the interaction between a LEGO excavator and the splashing waves. There are $334$,$815$ solid Gaussian kernels and $280$,$000$ fluid Gaussian kernels. Those volumetric Gaussians not only capture nonlinear dynamics of two-way coupled fluids and solids but can also be rasterized to realistically render with both diffuse and specular shading. GSP re-engineers several state-of-the-art techniques from neural surface reconstruction, specular-aware Gaussian shader, position-based tension, and AI inpainting to ensure the quality of both simulation and rendering with 3DGS.
1. Introduction
@@
Visualization and reconstruction of 3D scenes have been the core of 3D graphics and vision. Recent development of learning-based techniques such as the neural radiance fields (NeRFs) (Mildenhall et al., 2020) sheds new light on this topic. NeRF casts the reconstruction pipeline as a training procedure and delivers state-of-the-art results by encapsulating the color, texture, and geometry of the 3D scene into an implicit MLP net. Its superior convenience and efficacy inspire many follow-ups e.g., with improved visual quality (Liu et al., 2020), faster performance (Yu et al., 2021; Garbin et al., 2021), and sparser inputs (Jain et al., 2021; Yuan et al., 2022a). NeRF is designed for ray tracing. Image synthesis with NeRF has to follow the path integral, which is less suitable for real-time or interactive applications unless dedicated compression or acceleration methods are employed e.g., with NGP encoding (Müller et al., 2022). 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) provides an elegant alternative. As the name suggests, 3DGS learns a collection of Gaussian kernels from the input images. Apart from NeRF, a novel view of the scene from an unseen camera pose is generated using rasterization with the tile-splatting technique. Therefore, fast rendering with 3DGS is feasible.
@@
It is noted that Gaussian kernels not only serve as a good rendering agent but also explicitly encode rich information of the scene geometry. This feature suggests 3DGS a good candidate for dynamic scenes (Wu et al., 2023; Yang et al., 2023; Duisterhof et al., 2023), animated avatars (Moreau et al., 2023; Zielonka et al., 2023), or simulated physics (Xie et al., 2023). We expand on this intuition, enhancing the current 3DGS ecosystem by injecting physics-based fluid and solid interactions into a 3DGS scene. This appears straightforward at first sight. Since 3DGS kernels are essentially a collection of ellipsoids, they can be used for the discretization of the fluid and solid dynamics just as position-based dynamics (Macklin et al., 2016), oriented particles (Müller and Chentanez, 2011) or other particle-based simulation techniques. Unfortunately, a simple combination of those techniques does not yield the results expected. Large rotational deformation of the solid objects affects the splatting results with sharp and spiky noises. During fluid motion, fluid particles undergo substantial positional shifts, moving from the inside to the outside or vice versa. Fluids are both translucent and specular. The vanilla 3DGS simplifies the composition of the light field without well-defined appearance properties. This limitation makes fluid rendering cumbersome with 3DGS.
@@
This paper presents a system namely Gaussian Splashing (GSP), a 3DGS-based framework that enables realistic interactions between solid objects and fluids in a physically meaningful way, and thus generates two-way coupled fluids-solids dynamics in novel views. GSP integrates Lagrangian fluid and 3DGS scenes through a unified framework of position-based dynamics (PBD) (Müller et al., 2007; Macklin et al., 2016). We follow a recent contribution of Gaussian shader (Jiang et al., 2023) to augment Gaussian kernels with additional environmental information so that specular shading can be dynamically synthesized along with the fluid’s movement. For solid objects, GSP uses an anisotropy loss to cap the stretching ratio during 3DGS training and mitigate the rendering artifacts induced by rotational deformation. We approximate the normal of a fluid kernel based on the surface tension if it is near the fluid surface. For scattered fluid droplets, we resort to a depth volume rendered via the current camera pose to estimate the normal information (van der Laan et al., 2009). GSP is versatile, due to the flexibility of PBD. It handles deformable bodies, rigid objects, and fluid dynamics in a unified way. While it is possible to incorporate more complicated constitutional models as in (Feng et al., 2023) and (Li et al., 2023). We found that PBD-based simulation suffices in many situations. We further augment GSP with an image-space segmentation module to select objects of interest from the 3DGS scene. We exploit the latest generative AI to fill the missing pixels to enable interesting physics-based scene editing.
@@
Dynamic neural radiance fields generalize the original NeRF system to capture time-varying scenes e.g., by decomposing time-dependent neural fields into an inverse displacement field and canonical time-invariant neural fields (Park et al., 2021a, b; Tretschk et al., 2021; Weng et al., 2022), or estimating the time-continuous 3D motion field (Pumarola et al., 2021; Du et al., 2021; Li et al., 2021; Xian et al., 2021; Gao et al., 2021; Liu et al., 2022; Guo et al., 2023) with an added temporal dimension. Existing arts enable direct edits of NeRF reconstructions (Bao et al., 2023; Dong and Wang, 2023; Jiang et al., 2022). In dynamic scenes, the rendering process needs to trace deformed sample points back to rest-shape space to correctly retrieve the color/texture information (Yuan et al., 2022b; Qiao et al., 2023; Xu and Harada, 2022; Peng et al., 2022). They often extract a mesh/grid from the NeRF volume. It is also possible to integrate physical simulation with NeRF using meshless methods (Li et al., 2023; Feng et al., 2023). Point NeRF (Xu et al., 2022) and 3DGS (Kerbl et al., 2023) offer a different perspective to scene representation explicitly using points/Gaussian kernels to encode the scene. The success of 3DGS has inspired many studies to transplant techniques for dynamic NeRF to 3DGS (Wu et al., 2023; Liang et al., 2023b; Xu et al., 2023). They incorporate learning-based deformation and editing techniques to reconstruct or generate dynamics of NeRF scenes. It is noteworthy that a recent work from Xie et al. (2023) integrates physical simulation with 3DGS, leveraging the unified proxy for both simulating and rendering.
@@
Lagrangian fluid simulation tracks fluid motion using individual particles as they traverse the simulation domain. A seminal approach within this domain is smoothed particle hydrodynamics (SPH) (Monaghan, 1992), which solves fluid dynamics equations by assessing the influence of neighboring particles. Despite its efficacy, SPH, particularly in its standard and weakly compressible forms (WCSPH) (Becker and Teschner, 2007), suffers from parameter sensitivity, e.g., kernel radius and time-step size for stiff equations. To relax the time-step restriction, predictive-corrective incompressible SPH (PCISPH) (Solenthaler and Pajarola, 2009) iteratively corrects pressure based on the density error. Similarly, position-based dynamics (Müller et al., 2007) provides a robust method of solving a system of non-linear constraints using Gauss-Seidel iterations by updating particle positions directly, which can also be employed in fluid simulation (Macklin and Müller, 2013) with improved stability. Furthermore, surface tension can also be generated (Xing et al., 2022) using the position-based iterative solver by tracking surface particles and solving constraints on them to minimize the surface area.
@@
Achieving precise rendering of reflective surfaces relies on accurately estimating scene illumination, such as environmental light, and material properties like bidirectional reflectance distribution function (BRDF). This task falls under the domain of inverse rendering (Barron and Malik, 2014; Nimier-David et al., 2019). Some NeRF-related methodologies disentangle the visual appearance into lighting and material properties, which can jointly optimize environmental illumination, surface geometry, and material (Bi et al., 2020; Boss et al., 2021a, b; Srinivasan et al., 2021; Zhang et al., 2021a, b). Other NeRF studies (Liang et al., 2022, 2023a; Liu et al., 2023; Kang et al., 2021) aim to enhance the accuracy of the normal estimation in physically based rendering (PBR). Nevertheless, these efforts face challenges such as time-consuming training and slow rendering speed. On the contrary, 3DGS naturally offers a good normal estimation as the shortest axis of the Gaussian kernel (Guédon and Lepetit, 2023; Jiang et al., 2023). Following this idea, it is possible to achieve high-quality rendering of reflective objects and training efficiency simultaneously (Jiang et al., 2023).
@@
Point-based rendering has been an active topic in computer graphics since the 1980s (Levoy and Whitted, 1985). The simplest method (Grossman and Dally, 1998) renders a set of points with a fixed size. It suffers from holes and relies on post-processing steps such as hole-filling (Grossman and Dally, 1998; Pintus et al., 2011) to correct the resulting rendering. An improvement is to use ellipsoids instead of volume-less points. This strategy is usually referred to as splatting (Zwicker et al., 2001). The ellipsoids are rendered with a Gaussian alpha-mask to eliminate visible artifacts between neighboring splats and then combined by a normalizing blend function (Zwicker et al., 2001; Alexa et al., 2004). Point-based rendering well synergizes with Lagrangian fluid rendering, enabling the calculation of fluid thickness and depth through splatting. This approach (van der Laan et al., 2009) achieves fluid rendering at an impressive real-time speed. Further extensions to splatting aim to automatically compute the shape and color of ellipsoids, for example, auto splats (Childs et al., 2012). With the development of deep learning in recent years, learning-based approaches improve the image quality of splatting (Bui et al., 2018; Yang et al., 2020). 3DGS (Kerbl et al., 2023) has introduced this technique into 3D reconstruction, enabling high-quality real-time novel view synthesis for reconstructed scenes. A natural idea is to combine 3DGS with fluid rendering, enabling interaction between reconstructed scenes and dynamic fluids.
@@
Figure 2. An overview of GSP pipeline.  The input to our system comprises multi-view images that capture a 3D scene. During the preprocessing stage, foreground objects are isolated and reconstructed. This is followed by point sampling to facilitate scene discretization for PBD simulation and Gaussian rendering. We train the Gaussian kernels using differentiable 3DGS, which takes into account appearance materials and lighting conditions. These kernels are animated using PBD, in conjunction with fluid particles, to tackle the dynamics of both solids and fluids within the scene. Finally, the dynamic scene is rendered into images. This rendering process includes detailed modeling of specular reflections, thereby providing visually accurate representations of the simulated interactions between solids and fluids.
3. Background
@@
To make our paper self-contained, we start with a brief review of PBD and 3DGS on which our pipeline is built. More detailed discussions are available in the relevant literature e.g., (Macklin et al., 2016; Müller et al., 2007; Kerbl et al., 2023; Jiang et al., 2023).
3.1. Position-based Dynamics
@@
PBD/XPBD treats a dynamic system as a set of $N$ vertices and $M$ constraints. This perspective offers an easy and efficient simulation modality, converting the variational optimization to the so-called constraint projections. Specifically, XBPD considers the total system potential $U$ as a quadratic form of all the constraints $\bm{C}(\bm{x})=[C_{1}(\bm{x}),C_{2}(\bm{x}),...,C_{M}(\bm{x})]^{\top}$ such that $U=\frac{1}{2}\bm{C}^{\top}(\bm{x})\bm{\alpha}^{-1}\bm{C}(\bm{x})$. Here, $\bm{\alpha}$ is the compliance matrix i.e., the inverse of the constraint stiffness. XBPD estimates an update of constraint force (i.e., the multiplier) $\Delta\bm{\lambda}$ by solving:
(1) $\left[\Delta t^{2}\nabla\bm{C}(\bm{x})\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})+\bm{\alpha}\right]\Delta\bm{\lambda}=-\Delta t^{2}\bm{C}(\bm{x})-\bm{\alpha}\bm{\lambda},$

### v2

# Gaussian Splashing: Unified Particles for Versatile Motion Synthesis and Rendering

@@ Abstract
We demonstrate the feasibility of integrating physics-based animations of solids and fluids with 3D Gaussian Splatting (3DGS) to create novel effects in virtual scenes reconstructed using 3DGS. Leveraging the coherence of the Gaussian Splatting and Position-Based Dynamics (PBD) in the underlying representation, we manage rendering, view synthesis, and the dynamics of solids and fluids in a cohesive manner. Similar to GaussianShader, we enhance each Gaussian kernel with an added normal, aligning the kernel’s orientation with the surface normal to refine the PBD simulation. This approach effectively eliminates spiky noises that arise from rotational deformation in solids. It also allows us to integrate physically based rendering to augment the dynamic surface reflections on fluids. Consequently, our framework is capable of realistically reproducing surface highlights on dynamic fluids and facilitating interactions between scene objects and fluids from new views. For more information, please visit our project page at https://gaussiansplashing.github.io/.
@@
Figure 1: Versatile Motion Synthesis and Rendering.  Gaussian Splashing (GSP) is a unified framework combining 3D Gaussian Splatting (3DGS) and position-based dynamics. By leveraging their coherent point-based representations, GSP delivers high-quality rendering for novel dynamic views involving interacting deformable bodies, rigid objects, and fluids. GSP enables a variety of compelling effects and new human-computer interaction modalities not available with existing NeRF/3DGS-based systems. The teaser figure showcases a cliff battered by waves (a), a deformable ficus plant (b), flooding garden (c) and piled and scattered rigid lego bulldozers in a box (d). The reconstructed Gaussians not only capture the nonlinear dynamics of fluids and solids but can also be rasterized to realistically render with both diffuse and specular shadings. GSP re-engineers several state-of-the-art techniques from neural surface reconstruction, specular-aware Gaussian shading, position-based surface tension, and AI inpainting to ensure the quality of both simulation and rendering with 3DGS.
1 Introduction
@@
Visualization and reconstruction of 3D scenes have been the core of 3D graphics and vision. Recent development of learning-based techniques such as the neural radiance fields (NeRFs) [44] sheds new light on this topic. NeRF casts the reconstruction pipeline as a training procedure and delivers state-of-the-art results by encapsulating the color, texture, and geometry of the 3D scene into an implicit MLP net. Its superior convenience and efficacy inspire many follow-ups, e.g., with improved visual quality [38], faster performance [76, 19], and sparser inputs [25, 77]. NeRF is computationally expensive. Image synthesis with NeRF has to follow the path integral, which is less suitable for real-time or interactive applications unless dedicated compression or acceleration methods are employed e.g., with NGP encoding [50]. 3D Gaussian Splatting (3DGS) [29] provides an elegant alternative. As the name suggests, 3DGS learns a collection of Gaussian kernels from the input images.Apart from NeRF, a novel view of the scene from an unseen camera pose is generated using rasterization with the tile-splatting technique. Therefore, fast rendering with 3DGS is feasible.
@@
It is noted that Gaussian kernels not only serve as a good rendering agent but also explicitly encode rich information of the scene geometry. This feature suggests 3DGS a good candidate for dynamic scenes [67, 75, 16], animated avatars [46, 81], or simulated physics [69]. We expand on this intuition, enhancing the current 3DGS ecosystem by injecting physics-based fluid and solid interactions into a 3DGS scene. This appears straightforward at first sight. Since 3DGS kernels are essentially a collection of ellipsoids, they can be used for the discretization of the fluid and solid dynamics just as position-based dynamics [42], oriented particles [47] or other particle-based simulation techniques. Unfortunately, a simple combination of those techniques does not yield the results expected. Large rotational deformation of the solid objects affects the splatting results with sharp and spiky noises. During fluid motion, fluid particles undergo substantial positional shifts, moving from the inside to the outside or vice versa. Fluids are both translucent and specular. The vanilla 3DGS simplifies the composition of the light field without well-defined appearance properties. This limitation makes fluid rendering cumbersome with 3DGS.
@@
This paper presents a system namely Gaussian Splashing (GSP), a 3DGS-based framework that enables realistic interactions between solid objects and fluids in a physically meaningful way, and thus generates two-way coupled fluids-solids dynamics in novel views. GSP integrates Lagrangian fluid and 3DGS scenes through a unified framework of position-based dynamics (PBD) [48, 42]. We follow a recent contribution of GaussianShader [27] to augment Gaussian kernels with additional environmental information so that specular shading can be dynamically synthesized along with the fluid’s movement. For solid objects, GSP uses an anisotropy loss to cap the stretching ratio during 3DGS training and mitigate the rendering artifacts induced by rotational deformation. We approximate the normal of a fluid kernel based on the surface tension if it is near the fluid surface. For scattered fluid droplets, we resort to a depth volume rendered via the current camera pose to estimate the normal information [64]. GSP is versatile, due to the flexibility of PBD. It handles deformable bodies, rigid objects, and fluid dynamics in a unified way. While it is possible to incorporate more complicated constitutional models as in [17] and [31]. We found that PBD-based simulation suffices in many situations. We further augment GSP with an image-space segmentation module to select objects of interest from the 3DGS scene. We exploit the latest generative AI to fill the missing pixels to enable interesting physics-based scene editing.
@@
Dynamic neural radiance fields generalize the original NeRF system to capture time-varying scenes e.g., by decomposing time-dependent neural fields into an inverse displacement field and canonical time-invariant neural fields [52, 53, 63, 66], or estimating the time-continuous 3D motion field [56, 15, 33, 68, 18, 37, 22] with an added temporal dimension. Existing arts enable direct edits of NeRF reconstructions [3, 13, 26, 32]. In dynamic scenes, the rendering process needs to trace deformed sample points back to rest-shape space to correctly retrieve the color/texture information [78, 58, 72, 54]. They often extract a mesh/grid from the NeRF volume. It is also possible to integrate physical simulation with NeRF using meshless methods [31, 17]. Point NeRF [71] and 3DGS [29] offer a different perspective to scene representation explicitly using points/Gaussian kernels to encode the scene. The success of 3DGS has inspired many studies to transplant techniques for dynamic NeRF to 3DGS [67, 36, 73]. They incorporate learning-based deformation and editing techniques to reconstruct or generate dynamics of NeRF scenes. It is noteworthy that a recent work from Xie et al. [69] integrates physical simulation with 3DGS, leveraging the unified proxy for both simulating and rendering.
@@
Lagrangian fluid simulation tracks fluid motion using individual particles as they traverse the simulation domain. A seminal approach within this domain is smoothed particle hydrodynamics (SPH) [45], which solves fluid dynamics equations by assessing the influence of neighboring particles. Despite its efficacy, SPH, particularly in its standard and weakly compressible forms (WCSPH) [6], suffers from parameter sensitivity, e.g., kernel radius and time-step size for stiff equations. To relax the time-step restriction, predictive-corrective incompressible SPH (PCISPH) [59] iteratively corrects pressure based on the density error. Similarly, position-based dynamics [48] provides a robust method of solving a system of non-linear constraints using Gauss-Seidel iterations by updating particle positions directly, which can also be employed in fluid simulation [40] with improved stability. Furthermore, surface tension can also be generated [70] using the position-based iterative solver by tracking surface particles and solving constraints on them to minimize the surface area.
@@
Achieving precise rendering of reflective surfaces relies on accurately estimating scene illumination, such as environmental light, and material properties like bidirectional reflectance distribution function (BRDF). This task falls under the domain of inverse rendering [4, 51]. Some NeRF-related methodologies disentangle the visual appearance into lighting and material properties, which can jointly optimize environmental illumination, surface geometry, and material [7, 8, 9, 60, 79, 80]. Other NeRF studies [34, 35, 39, 28] aim to enhance the accuracy of the normal estimation in physically based rendering (PBR). Nevertheless, these efforts face challenges such as time-consuming training and slow rendering speed. On the contrary, 3DGS naturally offers a good normal estimation as the shortest axis of the Gaussian kernel [21, 27]. Following this idea, it is possible to achieve high-quality rendering of reflective objects and training efficiency simultaneously [27].
@@
Point-based rendering has been an active topic in computer graphics since the 1980s [30]. The simplest method [20] renders a set of points with a fixed size. It suffers from holes and relies on post-processing steps such as hole-filling [20, 55] to correct the resulting rendering. An improvement is to use ellipsoids instead of volume-less points. This strategy is usually referred to as point splatting [82]. The ellipsoids are rendered with a Gaussian alpha-mask to eliminate visible artifacts between neighboring splats and then combined by a normalizing blend function [82, 2]. Point-based rendering well synergizes with Lagrangian fluid rendering, enabling the calculation of fluid thickness and depth through splatting. This approach [64] achieves fluid rendering at an impressive real-time speed. Further extensions to splatting aim to automatically compute the shape and color of ellipsoids, for example, auto splats [12]. With the development of deep learning in recent years, learning-based approaches improve the image quality of splatting [11, 74]. 3DGS [29] has introduced this technique into 3D reconstruction, enabling high-quality real-time novel view synthesis for reconstructed scenes. A natural idea is to combine 3DGS with fluid rendering, enabling interaction between reconstructed scenes and dynamic fluids.
@@
Figure 2: An overview of GSP pipeline.  The input to our system comprises multi-view images that capture a 3D scene. During the preprocessing stage, foreground objects are isolated and reconstructed. This is followed by point sampling to facilitate scene discretization for PBD simulation and Gaussian rendering. We train the Gaussian kernels using differentiable 3DGS, which takes into account appearance materials and lighting conditions. These kernels are animated using PBD, in conjunction with fluid particles, to tackle the dynamics of both solids and fluids within the scene. Finally, the dynamic scene is rendered into images. This rendering process includes detailed modeling of specular reflections, thereby providing visually accurate representations of the simulated interactions between solids and fluids.
3 Background
@@
To make our paper self-contained, we start with a brief review of PBD and 3DGS on which our pipeline is built. More detailed discussions are available in the supplementary material and relevant literature e.g., [42, 48, 29, 27].
3.1 Position-Based Dynamics
@@
PBD/XPBD treats a dynamic system as a set of $N$ vertices and $M$ constraints. This perspective offers an easy and efficient simulation modality, converting the variational optimization to the so-called constraint projections. Specifically, XBPD considers the total system potential $U$ as a quadratic form of all the constraints $\bm{C}(\bm{x})=[C_{1}(\bm{x}),C_{2}(\bm{x}),...,C_{M}(\bm{x})]^{\top}$ such that $U=\frac{1}{2}\bm{C}^{\top}(\bm{x})\bm{\alpha}^{-1}\bm{C}(\bm{x})$. Here, $\bm{x}$ represents the position of vertices and $\bm{\alpha}$ is the compliance matrix, i.e., the inverse of the constraint stiffness. XBPD estimates an update of constraint force (i.e., the multiplier) $\Delta\bm{\lambda}$ by solving:
$\left[\Delta t^{2}\nabla\bm{C}(\bm{x})\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})+\bm{\alpha}\right]\Delta\bm{\lambda}=-\Delta t^{2}\bm{C}(\bm{x})-\bm{\alpha}\bm{\lambda},$ (1)